A Bibliography of Publications about Matched Field Processing

Christoph Mecklenbräuker

Signal Theory Group Dept. of Electrical Engineering

Ruhr-University Bochum D-44780 Germany

Tel: ?n/a? FAX: ?n/a?

E-mail: cfm@sth.ruhr-uni-bochum.de (Internet)

04 February 2012 Version 1.12

Abstract

This bibliography records publications about matched field processing.

Title word cross-reference

12-month [JCZ93].

24-month [JCZ94].

A.T.C [EJS82]. absorption [STW85]. Accuracy [ET86, Has84]. Acoustic [Opp84, CS92, CR92, HP94, Jen76, JF79, LO94a, LO94b, MLMS94, MN90, Por93, Sch95, Zio95, Bd76, E.84, Ger95, GG95a, Jes94]. Acoustics [BK93, JKPS94, Jenxx, AB79, IFKW78]. Adaptive [BS89, CZO87, Gab86, GZ94, Pre94, SSM95]. adjacent [BSE85]. Advanced [EJS82]. Algorithm [MB94c]. Algorithms [GG95a, Vib91, du 93, Ger94, Ger95]. Ambient [BS89, Cal64, LD85, Wil85]. Ambientes [Cla95]. Ambiguity [aDRM92]. Amplitude [Mar72, MN90]. Analyse [Mai89]. Analysis [Bri81, Opp84, Cal64, EJS81, MMB93b, PW93, Pri81a, Pri81b, Sle78, Tho82]. Analytical [LD85, Abr92]. Angel [EJS82]. Angle [EJS81]. Angular [EJS81]. Antenna [Gab86, Ma74]. Anwendung [Pur95]. Application [EJS82, Gra76, Joh82, Ma74, Thi87]. Applications [Efr75, PW93, RM71, Sto57, Tho80]. Approach [CS92, Pre94, TG95, Ger95, MN90, VS94]. Approximate [GMB95, KB90].

Approximation [Swi94]. Approximative [Kra93]. approximativen [Mai89]. Arguments [du 93]. ARMA [PG89]. Array [Böh91, BM94, Böh95a, Böh95b, GAB93, HC75, Hin77, KB90, KWG85, RG93, RS87, TG95, Vib89, Vib91, VO91, Has84, JCZ93, JCZ94, Jes94, JCxx, SS88, VS94]. Arrays [Ma74, VOK91, Abr92, E.84, Ham85]. Arrival [BK88, EJS81, EJS82, MB94c, RSZ94, Swi94]. assessment [BSE85]. Asymptotic [KB90, NL88]. Attenuation [KI77, Ing72]. Autocovariance [PG89]. average [Bd76]. Averaged [HK87].

Baltic [STW85]. Band [SM93, Smixx]. Based [VO91, CS89, CS92]. BAYESIAN [VS94]. Beamformer [SSM95]. Beamforming [CZO87, Xu94, Kro92]. beampower [E.84]. Bearing [Joh82]. below [Hol94]. Berücksichtigung [Kra87]. Bessel [Wat66, du 93]. Bochum [Böh89a, Böh83, Böh89b, Fel7x]. Boehme [Böh89a]. Bootstrap [Efr82, ET86, MB94a]. bottom [JCxx]. Bound [GMB95, NK94, aDRM91, Swi94]. boundaries [KI77]. Bounds [NV94]. Breaking [Hol94]. Breitbandverfahren [Mai95]. Broad [SM93]. Broad-Band [SM93]. Broadband [Jes93, MGB95, MGB96, Bd76]. bzw [Mai89].

C [FPJ93]. Calculating [BPW93]. Calculations [Bd76]. Calibration [RS87]. Case [Sle78]. channel [Kro92]. Channels [aDRM92]. Characteristics [LD85, Wil85]. Characterization [Tho77]. Classification [HP94]. Closely [Swi94]. closest [MN90]. coefficients [Ing72]. Coherence [SSM95]. coherent [KI77]. Coherently [HK87]. combination [Ger95]. Communications [BT58]. Comparison [BS89, NV94].

Complex [du 93]. component [KI77]. Components [EK94]. Composite [HC75]. Computation [du 93]. Computational [JKPS94]. computer [Bd76]. Computing [DW87]. Concepts [Vib89]. conditions [BSE85]. Confidence [ET86]. Constraint [Xu94]. continental [BSE85]. Continuous [Xu94]. Conventional [PW93]. correlated [BKS88]. Correlation [CS62, KI80]. correlator [SS88]. Coupled [FPJ93, Thi87]. coupled-mode [Thi87]. Covariance [HK87, Vib91]. Cramér [aDRM91, GMB95, NK94, Swi94]. Cramér-Rao [aDRM91, GMB95, NK94]. Criterion [MKBB91]. Current [De 79]. Curvature [Efr75]. CW [Jenxx].

Data [BM94, Böh95a, Böh95b, Bri81, GZ94, HC75, MB92, MMB93b, MB94a, Ger94, Ger95, GG95a, Jes94, JCxx]. Decomposition [DT84, HC75]. Defining [Efr75]. deformierbaren [Som62a]. Density [Cal64]. Dependent [LO94b]. Depth [Hin77, LS93]. Design [Hin77]. Detection [BM94, BS89, CMN94, GZ94, HP94, VOK91]. Deterministic [LS90]. Development [Tho77]. Differential [KP93, Øks85, RR93]. Differentialgleichungen [Som62b]. Diffraction [Dev84]. Digital [KKK81]. Direction [BK88, RSZ94, VS94]. Directions [Swi94]. Discrete [Sle78]. distributed [MN90]. distributions [Ger94]. DOA [NV94]. Dominant [HS92, HS93]. Dual [KDB93]. durch [Cla95].

Ebene [Pur95]. effect [Has84]. Effects [SSM95, Vib91, Jes91]. Efficiency [Efr75]. efficiently [AB79]. Elastizitätstheorie [LL53]. Electromagnetic [CR92]. Elektrotechnik [Böh83, Böh89a, Böh89b]. Elektrotechniker [Fel7x]. Empirical [KB90]. Engineering [BT58]. Environment [LS90, LO94a, LO94b, NL88, Pre94]. Environmental [FDR89, MGB95, BSE85, CK91, Jes91]. Environments [DW87, HS92, HS93, aDRM91]. equation [JK75]. Equations [KP93, Øks85, RR93]. Errors [ET86]. Estimated [PG89]. Estimates [KB90, Tho80]. Estimating [JCxx]. Estimation [BK88, BK91, EJS81, EJS82, Gab86, GMB95, GAB93, HK87, Joh82, KDB93, LO71, LTxx, MKBB91, MB92, Mar72, MT82, MGB95, NK94, NV94, PG89, PF83, RG93, SM93, Tho77, Tho82, VOK91, BKS88, E.84, Hin79]. Estimator [GMB95, NL88]. estimators [E.84]. Evaluation [du 93]. Experimental [GG95b, RG93]. Explicit [GMB95]. Explosive [Pek48]. Exponential [Efr78]. Extended [SS88].

F [MMB95, Pur95]. F-Test [MMB95]. F-Tests [Pur95]. Fachrichtung [Böh83, Böh89a, Böh89b]. Families [Efr78]. Far [RS87]. Far-Field [RS87]. Field [BK93, GMB95, KWG85, LO94a, LO94b, MB94c, MB94b, MMB95, MGB96, Pre94, RS87, TG95, Wil85, Zio95, Abr92, BKS88, FDR89, GG95a, Jes93, Jes94, Kro92]. Fields [BK88, Ros85, SVOK94]. Filtering [CMN94, HC75, Kai81, HS89]. Filters [GAB93, RG93]. FINDING [VS94]. First [du 93, IFKW78]. Fitting [Vib89, VOK91, VO91]. Flachwasser [Cla95, Pur95]. Florida [BSE85]. Fluids [MF53]. Focalization [CK91]. focusing [CK91]. Fourier [Sle78]. Frequencies [Hol94]. frequency [E.84, MN90]. frequency-wavenumber [E.84]. full [Jes94]. full-field [Jes94]. Function [NM65]. Functions [CS62, Roa70, Roa82, Sle78, Wat66, du 93]. Fundamentals [Zio95].

Gauss [Ger95]. Generalized [RM71]. generated [KI80]. Genetic [GG95a, Ger94, Ger95]. geo [Jes94]. geo-acoustic [Jes94]. Geoacoustic [FDR89, GG95b, JCxx]. geometric [GG95b]. Geometry [Efr78]. Geophysical [Dev84, JCZ93, JCZ94]. Global [GG95a]. Green [Roa70, Roa82].

Harmonic [Tho82]. heading [Has84]. hervorgerufen [Cla95]. High [EJS81]. Höhere [BHW85, MV90]. Horizontal [Hin77, Abr92, Ham85]. Hydroacoustic [GZ94]. Hydrodynamik [LL52]. Hz [Hol94].

Identification

[MMB93a, MLMS94, MB94c, MB94b].

Identifying [MMB95]. II

[HS93, LO94b, MF53, Tho77]. im [Pur95]. Image [AB79]. Implementierung [Cla95]. Inclusions [Xu94]. Independent [LO94a]. induced [Ham85]. Information [LTxx]. Ingenieure [BHW85]. Integer [du 93]. Integer-Order [du 93]. intensity [Bd76]. Interference [SSM95]. Intervals [ET86]. Introduction [Kat82, RR93, Sko62]. Inverse [BPW93, CR92, RM71, Tar86, Tho80]. Inversion [Ger94, Ger95, GG95b, GG95a, Jes94]. Issue [Imb94]. Iteration [BPW93]. IV [Smixx].

Jackknife [Efr82]. Jackknifing [Tho94]. Jacksonville [BSE85].

Kalman [GAB93, Kai81, RG93]. Kind [du 93]. Kleinste [Mai89]. Kleinste-Quadrate- [Mai89]. Known [HS92, SVOK94]. KRAKEN [Por91]. Kurzfassung [Böh83, Böh89a, Böh89b].

Large [DW87, Xu94]. Large-Scale [DW87]. layered [Bre60]. Least [BK88, KB90, LH74, MKBB91]. Lectures [Kai81]. Lehrbuch [Smixx]. Likelihood [KB90, Kra93, KDB93, Mai89, NL88, Hin79, HS89, Kra87]. line [Ham85]. Linear [Gra76, Kat66, Kat82, LO71]. Localization [CMN94, LS90, Abr92, BKS88, CK91, HS89, Jes91, SVOK94]. Location [BK91, DT84, KDB93, aDRM91, SM93]. Locations [RS87]. loss [FPJ93]. low [STW85]. Lower [Swi94].

Matched [BKS88, BK93, CMN94, Kro92, MB94c, MB94b, MMB95, MGB96, Pre94, Abr92, FDR89, Jes93]. Matched-field [Kro92, FDR89]. matching [Jes91]. Mathematical [Sto57]. Mathematik [BHW85, Fel7x, MV90, Smixx]. Matrices [RM71]. Matrix [DT84, HK87]. Maximum [Hin79, HS89, KB90, Kra93, KDB93, Mai89, NL88, Kra87]. Maximum-likelihood [Hin79, HS89]. Maximum-Likelihood-Schätzern [Kra87]. Maximum-Likelihood-Schätzung [Kra93]. Maximum-Likelihood-Zweistufenschätzers [Mai89]. means [Bd76]. Measured [Böh95b]. Measurement [BT58, LS93, STW85]. Measurements [Ing72]. Measures [ET86]. Measuring [Hin77]. Mechanik [Som62a]. Media [Bre60]. Medien [Som62a]. Medium [Che60]. Method [AB79, LD85, NM65, JK75]. Methoden [Böh83, Böh89a]. Methods [BK88, BK91, BR95, ET86, Joh82, MMB93b, MF53]. Microcomputers [DW87]. Minimax [Pre94]. Minimization [NM65]. minimum [Kro92]. mismatch [FDR89]. ML [GMB95, MGB95]. ML-Estimation [MGB95]. Mode [JF79, Xu94, FPJ93, HS89, Ing72, Jes91, Por91, Thi87]. mode-filtering [HS89]. Model [CS89, CS92, Gra76, JF79, MMB93a, PG89, Bd76, FPJ93, SR87]. Model-based [CS89, CS92]. Modeling [EK94, Jenxx, PG89]. Modelling [Jen76, JK75]. Models [CS62, GAB93, LO71, Por93, Sch95, Jen76]. Modenausbreitung [Pur95]. Modifikation [Kra87]. monitoring [JCZ93, JCZ94]. month [JCZ93, JCZ94]. Motion [MF53]. Moving [GZ94, LO94a, LO94b, RSZ94]. Multidimensional [HC75]. Multimode [MB94b, MMB95]. Multipath [HS92, HS93, LS90, LS93, LTxx, NL88, aDRM91, aDRM92]. Multiple [BM94, DT84, MB94a, MGB96, PF83, RSZ94, Tho94]. Multiple-Window [Tho94]. Multitaper [PW93]. Multitapers [BPW93].

Navigation [EJS82]. Near [KWG85]. Near-Field [KWG85]. Networks [HP94]. Neural [HP94]. Newton [Ger95]. Nichtparametrische [Böh83]. Noise [BK88, BS89, Cal64, CS62, GMB95, KWG85, LD85, Vib91, Wil85, BKS88, Ham85, KI80, SVOK94]. Nonstationary [MMB93a, MMB93b]. Nonwave [TG95]. Normal [JF79, Jes91, FPJ93, Por91]. Normal-Mode [JF79, Jes91]. Numerical [KP93].

Observations [Hol94]. Ocean

[BK93, Cal64, CS92, JKPS94, Jenxx,

MLMS94, MB94c, MB94b, MGB95, NK94, Bd76, Kro92, KI80]. Operators [Kat66, Kat82]. Optimal [SM93, SVOK94]. Optimum [SSM95, BKS88]. Optimum/Adaptive [SSM95]. Order [Efr75, du 93]. Ortung [Kra93, Mai95]. Other [Efr82, ET86]. overlapped [SS88].

Palaeoclimatology [Tho80]. parabolic [JK75]. Parameter [MB92, FDR89]. Parameters [HS92, HS93, HK87, LTxx, MKBB91, MGB95, NL88, PF83, GG95b]. Parametric [BK91, MMB93b, Vib91, VS94]. Parametrische [Böh89a]. Part [BT58, E.84, HS92, HS93, LO94a, LO94b, MF53, RS87, Tho77]. Partial [RR93].

PARTIALLY [SVOK94]. Partielle [Som62b]. Passive [CS89, CMN94, HS92, HS93, SR87, Has84, HS89]. pattern [E.84]. Performance [E.84, NV94, SSM95, SR87]. Perturbation [Kat66, Kat82]. phase [IFKW78]. Physical [PW93]. Physics [De 79, MF53]. Physik [Som62b]. placement [Has84]. Plans [Efr82]. Point [BT58, LO94a, LO94b, MN90]. position [Hin79]. posteriori} [Ger94]. Power [BT58, GMB95]. pre [BSE85]. pre-assessment [BSE85]. Predict [LD85]. Presence [BK88, GMB95]. Prewhitening [MGB96]. Principles [Uri67]. Probabilistic [MLMS94]. Probability [Cal64]. Problem [Efr75, Tar86, BKS88]. Problems [Joh82, LH74]. Processing [BK93, Böh91, Böh95b, CS92, EJS82, Gab86, KKK81, KB90, MB94c, MB94b, MMB95, MGB96, Pre94, TG95, Vib89, Vib91, VO91, Zio95, Abr92, BKS88, FDR89, Jes93]. processor [SS88]. program [Por91]. Programs [du 93]. Progress [JCZ93, JCZ94]. Prolate [Sle78]. Propagation [Che60, Jen76, JF79, Jenxx, KWG85, LS93, MB94b, MMB95, Pek48, Pre94, Sob85, FPJ93, JK75, KI77, Thi87]. properties [JCxx]. Pulse [Jenxx].

Quadrate [Mai89]. Quadratic [Tho80]. Quadratic-Inverse [Tho80].

Radar [Sko62]. radiating [Hin79]. Radio [Chr85, Imb94]. RAIBAC [Bd76]. Random [Che60, LS90, NK94, Ros85, E.84, Kro92]. randomly [Thi87]. Range [LO94a, LO94b, NK94]. Range-Dependent [LO94b]. Range-Independent [LO94a]. Ranging [CS89, HS92, HS93, Has84, SR87]. Rao [GMB95, NK94, aDRM91, Swi94]. Rauschen [Cla95]. Ray [MLMS94]. Receivers [BS89]. Remote [LS93]. Report [JCZ93, JCZ94, IFKW78]. Resampling [Efr82]. Resistant [MT82]. Resolution [EJS81]. Response [KWG85]. responses [Ham85]. Results [KB90, RG93, GG95b]. Retrieving [Böh95a]. Reverberation [EK94, Bd76]. Robust [BS89, CZO87, Mar72, MT82, MGB96, PG89]. Robust-Resistant [MT82]. room [AB79]. rough [KI77]. Ruhr [Böh83, Böh89a, Böh89b, Fel7x]. Ruhr-Universität [Böh83, Böh89a, Böh89b, Fel7x].

SACLANTCEN [FPJ93, JF79, Sch95]. salinity [STW85]. Scale [DW87]. Scattering [CR92, EJS81]. Schätzern [Kra87]. Schätzung [Kra93]. Sea [GZ94, Hol94, STW85]. seafloor [JCZ93, JCZ94]. Second [Efr75, du 93]. Section [MF53]. Seismic [Böh95b, MMB93a, MMB93b, MB94a]. Seismik [Mai95]. seismoacoustic [Ger94]. sensitivity [Jes94]. Sensor [KB90, TG95, Vib89, VOK91, VO91]. Sensorgruppen [Mai95, Kra93]. sensors [MN90]. Sequences [Ros85]. Series [Bri81, Pri81a, Pri81b]. several [Jen76]. Shallow [GZ94, IFKW78, MB94c, MB94b, MGB95, Pek48, Abr92, FDR89, GG95a, GG95b, Ham85, Ing72, Jes91, Jes93, JCZ93, JCZ94, Jes94, KI77, Thi87]. Shallow-Water [IFKW78, FDR89]. Shape [RS87]. shelf [BSE85]. Ship [Cal64, KWG85]. Shipping [LD85]. Short [Kat82]. Signal [BM94, Böh95b, BS89, CS92, EJS82, GZ94, GMB95, KKK81, MKBB91, Mar72, MB94c, Zio95]. Signale [Böh93]. Signalentdeckung [Mai95]. Signals [Böh95a, LS90, MMB93a, Bd76, Jes93, SVOK94]. Signalschätzung [Kra93]. Simple [GMB95, Swi94]. Simplex [NM65]. simulating [AB79]. Skript [Fel7x]. small [AB79]. small-room [AB79]. SNAP [FPJ93, JF79, SR87]. Solution [KP93]. Solving [LH74]. some [du 93]. Sonar [Böh95b, CMN94, KKK81, Por93]. Sonaranwendungen [Mai95]. Sonardaten [Pur95]. Sound [Hol94, Pek48, Uri67, JK75, KI77, STW85, Thi87]. Source [BKS88, BK91, DT84, Hin77, KDB93, LS90, LTxx, LO94a, LO94b, NK94, NL88, SM93, Abr92, CK91, Hin79]. Sources [HK87, LS93, MGB95, PF83, RS87, Swi94]. Space [Zio95]. Space-Time [Zio95]. Spaced [Swi94]. Spacial [PF83]. Spatial [CS62, KI80]. Spatial-Correlation [CS62]. spatially [MN90]. Special [Imb94]. Spectra [BT58, Tho94]. Spectral [BPW93, Gab86, Joh82, PG89, PW93, PF83, Pri81a, Pri81b]. Spectrum [EJS81, MT82, Tho77, Tho80, Tho82, E.84]. speed [Thi87]. Spektralanalyse [Böh83, Böh89b]. Spheroidal [Sle78]. Squares [BK88, KB90, LH74, MKBB91]. SSK [Böh89a]. Standard [ET86]. stationärer [Mai89]. Stationary [Ros85]. Statistical [Böh95b, Efr75, ET86, HK87, LD85]. Stochastic [BR95, KP93, Øks85, Sob85]. Stochastische [Böh93]. stratified [KI80]. Structure [aDRM92, Xu94]. study [Abr92, Jes94, SR87]. sub [Has84]. sub-array [Has84]. Suboptimal [SM93]. Subspace [Vib89, VOK91, VO91]. Sufficiency [HK87]. summary [IFKW78]. Suppression [EK94]. surface [KI80]. Surveillance [EJS82]. System [Opp84, Jes91]. Systemidentifikation [Böh89a, Böh89b]. Systems [EJS82, Gab86, Por93, aDRM91, Sko62].

Targets [RSZ94]. Techniques [EJS82, MGB96, PW93, Tho77]. Telescopes [Chr85, Imb94]. Teory [Sto57]. Terrain [EJS81]. Test [MMB95]. Testing [MB94a]. Tests [Pur95]. Theoretical [GAB93, MF53, Ham85]. Theorie [Cla95]. Theory [Bri81, Gra76, Kat66, Kat82, Ma74, Pek48, Tar86, Wat66, Zio95, Thi87]. Thomson [BPW93]. Time [Bri81, Pri81a, Pri81b, Zio95, MN90].

Tomography [Dev84, MLMS94]. Tool [MLMS94]. Topics [De 79]. Towed [BM94, GAB93, KWG85, RG93, JCZ93, JCZ94, Jes94, JCxx, SS88]. towed-array [SS88]. Tracking [BM94, RSZ94, MN90]. Transient [MB92, Jes93]. Transients [HP94]. Treatise [Wat66].

Umgebungsrauschen [Kra87]. Uncertain [Pre94]. Uncertainty [Sle78, Has84, VS94]. Uncertainty- [Sle78]. Underwater [BS89, Opp84, HP94, Hol94, LTxx, LO94a, LO94b, Uri67]. Univariate [PW93]. Universität [Böh83, Böh89a, Böh89b, Fel7x]. Unknown [BK88, GMB95, HS93, MGB95, RS87, Vib91, Xu94, SVOK94]. Untersuchung [Mai89]. Use [LTxx, JK75]. Using [Gab86, GZ94, GAB93, HP94, LS90, LS93, MGB95, RG93, RS87, VOK91, E.84, Ger94, Ger95, GG95a, HS89, MB94a, MB94c, MGB96].

V [Sle78]. Validation [MMB93a].

variability [FDR89]. variance [Kro92]. Various [CS62]. varying [Thi87]. Velocity [CMN94]. Verfahren [Kra93]. vertical [Ham85]. verwandte [Kra93]. via [HC75]. View [BT58]. Vorlesung [Böh83, Böh89a, Böh89b, Fel7x].

Water

[IFKW78, Pek48, Sto57, Abr92, FDR89, GG95a, GG95b, Ham85, Ing72, Jes91, Jes93, JCZ93, JCZ94, Jes94, KI77, STW85, Thi87]. Wave [Che60, MB92, Sle78, Sob85, Xu94]. Wavefields [MKBB91]. Wavefront [HC75]. Waveguide [Hin77, NK94, Tho77, Xu94, Hin79]. wavenumber [E.84]. Waves [Bre60, Hol94, Sto57]. Weighted [MKBB91, VOK91]. Wellen [Pur95]. Wellenfelder [Mai89]. Wellenmodell [Cla95]. Wellenparameter [Kra87]. Wideband [BM94, HK87, KDB93]. Wiener [Kai81]. wind [Ham85, Cla95]. wind-induced [Ham85]. Window [MGB96, Tho94]. WT [Tho77].

zur [Böh89a, Kra93, Mai89, Mai95]. Zweistufenschätzers [Mai89].

References

Allen:1979:IME

[AB79] J. B. Allen and D. A. Berkley.

Image method for efficiently simulating small-room acoustics. J. Acoust. Soc. Am., 65(4):943–950, April 1979.

Abraham:1992:ASS

[Abr92] D. Abraham. An analytical study of source localization in shallow water by matched–field processing of horizontal arrays. SACLANT-CEN Memorandum SM-255, La Spezia, August 1992.

Rendas:1991:CRB

[aDRM91] M. Jo ao D. Rendas and José M. F.

Moura. Cramér-rao bound for location systems in multipath environments. IEEE Trans. Signal Processing, 39(12):2593–2610, December 1991.

Rendas:1992:ASM

[aDRM92] M. Jo ao D. Rendas and José M. F. Moura. Ambiguity structure of multipath channels. In Proc. IEEE ICASSP, pages 509–512. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992.

Bauer:1859:xxx

[Bau59] Bauer. Journal für Mathematik, LVI:104–106, 1859.

Bachmann:1976:CRA

[Bd76] W. Bachmann and B. de Raigniac. Calculations of the reverberation and average intensity of broadband acoustic signals in the ocean by means of the RAIBAC computer model. J. Acoust. Soc. Am., 59:31–38, 1976.

Burg:1985:HMF

[BHW85] Burg, Haf, and Wille. Höhere Mathematik für Ingenieure. Teubner-Verlag, Stuttgart, Germany, 1985?

Boehme:1988:LSM

[BK88] J. F. Böhme and D. Kraus. On least squares methods for direction of arrival estimation in the presence of unknown noise fields. In Proc. IEEE ICASSP, pages 2833–2836. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 11-14 1988.

Boehme:1991:PMS

[BK91] J. F. Böhme and D. Kraus. Parametric methods for source location estimation. In 9^{th} IFAC/IFORS Symp. on Identification and System Parameter Estimation. Budapest, 1991.

Baggeroer:1993:MFP

[BK93] A. B. Baggeroer and W. A. Kuperman. Matched field processing in ocean acoustics. In I. M. G.

Lourtie and J. M. F. Moura, editors, Acoustic Signal Processing for Ocean Exploration,NATO ASI Series C, vol. 388, pages 79–114. Kluwer Academic Publisher, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993.

Baggeroer:1988:MFP

[BKS88] A. B. Baggeroer, W. A. Kuperman, and H. Schmidt. Matched field processing: Source localization in correlated noise as an optimum estimation problem. J. Acoust. Soc. Am., 83(2):571–587, February 1988.

Boehme:1994:MWS

[BM94] J. F. Böhme and D. Maiwald.

Multiple wideband signal detection and tracking from towed array data. In M. Blanke and T. Söderström, editors, SYSID 94, 10th IFAC Symposium on System Identification, volume 1, pages 107–112. Danish Automation Soc., Copenhagen, July 4-6 1994.

Boehme:1983:SNM

[Böh83] J. F. Böhme. Spektralanalyse, Nichtparametrische Methoden. Kurzfassung einer Vorlesung der Fachrichtung Elektrotechnik an der Ruhr-Universität Bochum, July 1983. bearbeitet von R. Knötsch.

Boehme:1989:PMS

[Böh89a] J. F. Böhme. Parametrische methoden zur boehme:1989:SSK und systemidentifikation, kurzfassung einer vorlesung der fachrichtung elektrotechnik an der ruhruniversität bochum, 1989. bearbeitet von A. Zoubir.

Boehme:1989:SSK

[Böh89b] J. F. Böhme. Spektralanalyse und Systemidentifikation, Kurzfassung einer Vorlesung der Fachrichtung Elektrotechnik an der Ruhr-Universität Bochum, 1989.

Boehme:1991:AP

[Böh91] J. F. Böhme. Array processing. In S. Haykin, editor, Advances in Spectrum Analysis and Array Processing, pages 1–63. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991.

Boehme:1993:SS

[Böh93] J. F. Böhme. Stochastische Signale. Teubner-Verlag, Stuttgart, Germany, 1993.

Boehme:1995:RSA

[Böh95a] J. F. Böhme. Retrieving signals from array data. In E. S. Husebye, editor, Selected papers from NATO ASI on Monitoring a Comprehensive Test Ban Treaty, pages 1–24. Kluwer Academic Publisher, Norwell, MA, USA, and Dordrecht, The Netherlands, January 23 – February 2 1995.

Boehme:1995:SAS

[Böh95b] J. F. Böhme. Statistical array signal processing of measured sonar and seismic data. In Proc. SPIE 2563 Advanced Signal Processing Algorithms, page (to appear). San Diego, July 1995.

Bell:1993:CTS

[BPW93] B. Bell, D. B. Percival, and A. T.

Walden. Calculating Thomson’s spectral multitapers by inverse iteration. J. Comput. and Graph. Stat., 1993.

Boender:1995:SM

[BR95] C. G. E. Boender and H. E.

Romeijn. Stochastic methods. In R. Horst and P. M. Pardalos, editors, Global Optimization, volume 2 of Nonconvex Optimization and Its Applications, pages 829–869. Kluwer Academic Publisher, Norwell, MA, USA, and Dordrecht, The Netherlands, 1 edition, 1995. ISBN 0-7923-3120-6, Bibl: M3 Hor 95 70.

Brekhovskikh:1960:WLM

[Bre60] L. M. Brekhovskikh. Waves in layered Media. Academic Press, New York, NY, USA, 1960. LCCN QC231 .B85 1960.

Brillinger:1981:TSD

[Bri81] D. R. Brillinger. Time Series:

Data Analysis and Theory. International Series in Decision Processes. Holt, Rinehart, and Winston, New York, NY, USA, 1981. ISBN 0-03-076975-2. various pp. LCCN QA280 .B74.

Bouvet:1989:CAR

[BS89] M. Bouvet and S. C. Schwartz.

Comparison of adaptive and robust receivers for signal detection in ambient underwater noise. IEEE Trans. ASSP, ASSP-37(4): 621, 1989.

Bucca:1985:PAE

[BSE85] P. J. Bucca, C. E. Sellinger, and Matthews J. E. A pre-assessment of environmental conditions on the continental shelf adjacent to Jacksonville, Florida. Technical report, NORDA, 1985.

Blackman:1958:MPS

[BT58] R. B. Blackman and J. W. Tukey. The measurement of power spectra from the point of view of communications engineering – part I. The Bell System Technical Journal, pages 185–282, 1958. CODEN BSTJAN. ISSN 0005-8580.

Calderon:1964:PDA

[Cal64] M. A. Calderon. Probability density analysis of ocean ambient and ship noise. Technical report, U.S. Navy Electron. Lab. Rep. 1248, November 1964.

Chernov:1960:WPR

[Che60] L. A. Chernov. Wave Propagation in a Random Medium. Dover Publications, New York, 1960. 168 pp. LCCN QC231 .C453.

Chow:1975:xxx

[Cho75] P. L. Chow. SIAM Rev., 17:57–81, 1975.

Christiansen:1985:RT

[Chr85] E. J. Christiansen. Radio Telescopes. Cambridge University Press, Cambridge, UK, 1985.

Collins:1991:FEF

[CK91] M. D. Collins and W. A. Kuperman. Focalization: environmental focusing and source localization.

J. Acoust. Soc. Am., 90(3):1410– 1422, 1991.

Clasen:1995:ARF

[Cla95] M. Clasen. Ambientes Rauschen in Flachwasser hervorgerufen durch Wind: Das Wellenmodell, Theorie und Implementierung. Technical report, Ruhr-Universität Bochum, Bochum, Germany, 1995.

Chan:1994:PSD

[CMN94] Y. T. Chan, S. P. Morton, and G. H. Niezgoda. Passive sonar detection and localization by matched velocity filtering. In Seventh SP Workshop on Statistical Signal & Array Processing, pages 337–340. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 26-29 1994.

Colton:1992:IAE

[CR92] D. Colton and Kress R. Inverse Acoustic and Electromagnetic Scattering. Applied Mathematical Sciences 93. Springer-Ver-lag, 1992. ISBN 0-387-55518-8, 3-540-55518-8, Bibl: M3 93 74.

Cron:1962:SCF

[CS62] B. F. Cron and C. H. Sherman. Spatial-correlation functions for various noise models. J. Acoust. Soc. Am., 34(11):1732– 1736, November 1962.

Candy:1989:MBP

[CS89] J. V. Candy and E. J. Sullivan.

Model-based passive ranging. J.

Acoust. Soc. Am., 85(6), June 1989.

Candy:1992:OAS

[CS92] J. V. Candy and E. J. Sullivan. Ocean acoustic signal processing: A model-based approach. J. Acoust. Soc. Am., 92(6), December 1992.

Cox:1987:RAB

[CZO87] H. Cox, R. M. Zeskind, and M. M. Owen. Robust adaptive beamforming. IEEE Trans. ASSP, ASSP-35(10):1365, 1987.

DeSanto:1979:TCP

[De 79] J. A. De Santo, editor. Topics in Current Physics, volume 8. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979.

Devaney:1984:GDT

[Dev84] A. J. Devaney. Geophysical diffraction tomography. IEEE Trans. Geos. Remote Sensing, GE-22(1), 1984.

Di:1984:MDM

[DT84] A. Di and L. Tian. Matrix decomposition and multiple source location. Proc. IEEE ICASSP, pages Sec. 33–4, 1984.

duToit:1993:ESA

[du 93] C. F. du Toit. Evaluation of some algorithms and programs for the computation of integer-order bessel functions of the first and second kind with complex arguments. IEEE Antennas and Propagation Magazine, pages 19–25, June 1993.

Dennis:1987:NCE

[DW87] J. E. Dennis, Jr. and D. J.

Woods. New computing environments: Microcomputers in largescale computing. In A. Wouk, editor, New Computing Environments: Microcomputers in Large-Scale Computing, pages 116–122. SIAM, Philadelphia, PA, USA, 1987.

E:1984:EFW

[E.84] Masry E. The estimation of the frequency-wavenumber spectrum using random acoustic arrays – part I. performance of beampower pattern estimators. J. Acoust. Soc. Am., 76(1):139–149, July 1984.

Efron:1975:DCS

[Efr75] B. Efron. Defining the curvature of a statistical problem (with applications to second order efficiency). Ann. Statist., 3(6):1189– 1242, 1975.

Efron:1978:GEF

[Efr78] B. Efron. The geometry of exponential families. Ann. Statist., 6 (2):362–376, 1978.

Efron:1982:JBO

[Efr82] B. Efron. The Jackknife, the Bootstrap, and Other Resampling Plans. SIAM, Philadelphia, PA, USA, 1982. ISBN 0-89871-179-7. vii + 92 pp. LCCN QA276.8 .E375 1982.

Evans:1981:HRA

[EJS81] J. F. Evans, J. R. Johnson, and D. F. Sun. High resolution angular spectrum estimation terrain scattering analysis and angle of arrival estimation. Proc. on 1^{st} ASSP Workshop on Spectral Estimation and cento, page 134, 1981.

Evans:1982:AAS

[EJS82] J. F. Evans, J. R. Johnson, and D. F. Sun. Application of advanced signal processing techniques to angel of arrival estimation in A.T.C. navigation and surveillance systems. Technical Report 582, MIT Lincoln Lab, Lexington, MA, 1982.

Edelson:1994:MSR

[EK94] G. S. Edelson and I. P. Kirsteins. Modeling and suppression of reverberation components. In Seventh SP Workshop on Statistical Signal & Array Processing, pages 437–440. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 26-29 1994.

Efron:1986:BMS

[ET86] B. Efron and R. Tibshirani. Bootstrap methods for standard errors, confidence intervals, and other measures of statistical accuracy. Statistical Science, 1(1):54– 77, 1986.

Feuillade:1989:EMS

[FDR89] C. Feuillade, D. R. Del Balzo, and M. M. Rowe. Environmental mismatch in shallow-water matchedfield processing: Geoacoustic parameter variability. J. Acoust. Soc. Am., 85(6), June 1989.

Felbecker:19xx:MFE

[Fel7x] Felbecker. Mathematik für Elektrotechniker, Skript einer Vorlesung an der Ruhr-Universität Bochum, 197x.

Ferla:1993:CSC

[FPJ93] C. M. Ferla, M. B. Porter, and F. B. Jensen. C–SNAP: Coupled SACLANTCEN normal mode propagation loss model. SACLANTCEN Memorandum SM-274, La Spezia, December 1993.

Gabriel:1986:USE

[Gab86] W. F. Gabriel. Using the spectral estimation in adaptive processing antenna systems. IEEE Trans. Antennas and Propagation, 34(3): 291–300, March 1986.

Gray:1993:TAE

[GAB93] D. A. Gray, B. D. O. Anderson, and R. R. Bitmead. Towed array estimation using kalman filters – theoretical models. IEEE J. Oceanic Eng., 18(4):543–556, October 1993.

Gegenbauer:1874:xx

[Geg74] Gegenbauer. Wiener Sitzungsberichte, LXX(2):433ff, 1874.

Gerstoft:1994:ISD

[Ger94] P. Gerstoft. Inversion of seismoacoustic data using genetic algorithms and a posteriori distributions. J. Acoust. Soc. Am., 95(2): 770–782, February 1994.

Gerstoft:1995:IAD

[Ger95] P. Gerstoft. Inversion of acoustic data using a combination of genetic algorithms and the gauss– newton approach. J. Acoust. Soc. Am., 97(4):2181–2190, April 1995.

Gingras:1995:GIA

[GG95a] D. F. Gingras and P. Gerstoft. Global inversion of acoustic field data in shallow water using genetic algorithms. In H. Schmidt, O. Diachok, A. Caiti, and P. Gerstoft, editors, Proc. of Full-field inversion methods in ocean and seismo-acoustics. Kluwer Academic Publisher, Norwell, MA, USA, and Dordrecht, The Netherlands, Spring 1995.

Gingras:1995:IGG

[GG95b] D. F. Gingras and P. Gerstoft. Inversion for geometric and geoacoustic parameters in shallow water: Experimental results. J. Acoust. Soc. Am., 97(6):3589– 3598, June 1995.

Gershman:1995:MES

[GMB95] A. B. Gershman, A. L. Matveyev, and J. F. Böhme. ML estimation of signal power in the presence of unknown noise field – simple approximate estimator and explicit cramer-rao bound. In Proc. IEEE ICASSP. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 4-8 1995.

Graybill:1976:TAL

[Gra76] F. A. Graybill. Theory and Application of the Linear Model.

Wadsworth Publishing Inc., Pacific Grove, California, 1976. ISBN 0-87872-108-8. xiv + 704 pp. LCCN QA279 .G7.

Gershman:1994:ADM

[GZ94] A. B. Gershman and V. A. Zverev.

Adaptive detection of moving signal using shallow sea hydroacoustic data. In Proc. IEEE ICASSP, pages II–297–II–300. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 19-23 1994.

Hamson:1985:TRV

[Ham85] Rachel M. Hamson. The theoretical responses of vertical and horizontal line arrays to windinduced noise in shallow water. J. Acoust. Soc. Am., 78(5):1702– 1712, November 1985.

Hassab:1984:EUH

[Has84] J. C. Hassab. The effect of uncertainty in the heading or placement of a sub-array on passive ranging accuracy. J. Acoust. Soc. Am., 75 (2):479–485, February 1984.

Halpeny:1975:CWD

[HC75] O. Halpeny and D. G. Childers.

Composite wavefront decomposition via multidimensional filtering of array data. IEEE Trans. Circuits ans Systems, 22(6), June 1975.

Hinich:1977:ADM

[Hin77] M. J. Hinich. Array design for measuring source depth in a horizontal waveguide. SIAM J. of Applied Math., 32(4):800–806, July 1977.

Hinich:1979:MLE

[Hin79] M. J. Hinich. Maximum-likelihood estimation of the position of a radiating source in a waveguide. J. Acoust. Soc. Am., 66(2):480–483, August 1979.

Hung:1987:SSC

[HK87] H. Hung and M. Kaveh. On the statistical sufficiency of the coherently averaged covariance matrix for the estimation of the parameters of wideband sources. In Proc. IEEE ICASSP, volume 1, pages 33–36. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 1987.

Hollett:1994:OUS

[Hol94] R. D. Hollett. Observations of underwater sound at frequencies below 1500 hz from breaking waves at sea. J. Acoust. Soc. Am., 95(1): 165–170, January 1994.

Hemminger94a

[HP94] T. L. Hemminger and Y. H. Pao. Detection and classification of underwater acoustic transients using neural networks. IEEE Trans Neural Networks, 5(5):712–718, September 1994.

Hinich:1989:MLP

[HS89] M. J. Hinich and E. J. Sullivan. Maximum-likelihood passive localization using mode-filtering. J. Acoust. Soc. Am., 85(1):214– 219, January 1989.

Hamilton:1992:PRM

[HS92] M. Hamilton and P. M. Schultheiss.

Passive ranging in multipath dominant environments, part I: Known multipath parameters. IEEE Trans. Signal Processing, 40(1):1– 12, January 1992. See also [HS93].

Hamilton:1993:PRM

[HS93] M. Hamilton and P. M. Schultheiss.

Passive ranging in multipath dominant environments, part II: Unknown multipath parameters. IEEE Trans. Signal Processing, 41 (1):1–12, January 1993.

Ingenito:1978:SWA

[IFKW78] F. Ingenito, R. Ferris, W. A. Kuperman, and S. N. Wolf. Shallowwater acoustics, summary report (first phase). Technical Report NRL Report 8179, U.S. Naval Research Laboratory, Washington, DC, 1978.

Imbriale:1994:SIR

[Imb94] Special issue on radio telescopes.

In W. A. Imbriale and M. Thorburn, editors, Proceedings of the IEEE, volume 82, pages 629–828. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1994.

Ingenito:1972:MMA

[Ing72] F. Ingenito. Measurements of mode attenuation coefficients in shallow water. J. Acoust. Soc. Am., 53(3):858–862, July 1972.

Jesus:19xx:EGB

[JCxx] S. M. Jesus and A. Caiti. Estimating geoacoustic bottom properties from towed array data. J. Computational Acoust., 19xx. Submitted.

Jesus:1993:GSM

[JCZ93] S. M. Jesus, A. Caiti, and H. Zambujo. Geophysical seafloor monitoring with a towed array in shallow water, 12-month progress report. Technical Report MAS2-CT920022, Economic European Community, DG XII, December 1993.

Jesus:1994:GSM

[JCZ94] S. M. Jesus, A. Caiti, and H. Zambujo. Geophysical seafloor monitoring with a towed array in shallow water, 24-month progress report. Technical Report MAS2-CT920022, Economic European Community, DG XII, December 1994.

Jensen:1976:APW

[Jen76] F. B. Jensen. Acoustic propagation modelling: why several models? Technical Report SACLANT-CEN CP-19, SACLANT ASW Research Centre, La Spezia/Italy, 1976. AD CO 10187.

Jensen:19xx:CPP

[Jenxx] F. B. Jensen. CW and pulse propagation modeling on ocean acoustics. 19xx.

Jesus:1991:NMM

[Jes91] S. M. Jesus. Normal-mode matching localization in shallow water: environmental and system effects. J. Acoust. Soc. Am., 90: 2034–2041, 1991.

Jesus:1993:BMF

[Jes93] S. M. Jesus. Broadband matched field processing of transient signals in shallow water. J. Acoust. Soc. Am., 93(4):1841–1850, April 1993.

Jesus:1994:SSF

[Jes94] S. M. Jesus. A sensitivity study for full-field inversion of geo-acoustic data with a towed array in shallow water. In L. Bjørnø, editor, Proc. 2^{nd} Europ. Conf. Underwater Acoustics, volume 2, pages 899–904. European Commission, Copenhagen, July 1994.

Jensen:1979:SSN

[JF79] F. B. Jensen and M. C. Ferla.

SNAP: The SACLANTCEN Normal-Mode Acoustic Propagation Model. SACLANT ASW Research Centre, La Spezia/Italy, January 1979.

Jensen:1975:UPE

[JK75] F. B. Jensen and H. R. Krol.

The use of the parabolic equation method in sound propagation modelling. Technical Report SACLANTCEN SM-72, SACLANT ASW Research Centre, La Spezia/Italy, 1975.

Jensen:1994:COA

[JKPS94] F. B. Jensen, W. A. Kuperman, M. B. Porter, and H. Schmidt. Computational Ocean Acoustics. American Institute of Physics Press, New York, 1994. ISBN 1-56396-209-8. xvi + 612 pp. LCCN QC242.2 .C66 1994.

Johnson:1982:ASE

[Joh82] D. Johnson. Application of spectral estimation methods for bearing estimation problems. Proc IEEE, 70(6):1018–1028, September 1982.

Kailath:1981:LWK

[Kai81] T. Kailath. Lectures on Wiener and Kalman Filtering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981.

Kato:1966:PTL

[Kat66] Tosio Kato. Perturbation Theory for Linear Operators. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. xix + 592 pp. LCCN QA320 .K33.

Kato:1982:SIP

[Kat82] Tosio Kato. A Short Introduction to Perturbation Theory for Linear Operators. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90666-5. xiii + 161 pp. LCCN QA329.2 .K3725 1982.

Kraus:1990:AER

[KB90] D. Kraus and J. F. Böhme. Asymptotic and empirical results on approximate maximum likelihood and least squares estimates for sensor array processing. In Proc. IEEE ICASSP, pages 1379– 1384. Albuquerque, 1990.

Kraus:1993:EDM

[KDB93] D. Kraus, A. Dhaouadi, and J. F.

Böhme. EM dual maximum likelihood estimation for wideband source location. In Proc. IEEE ICASSP, pages I–257, I–260. Minneapolis, 1993.

Kuperman:1977:ACC

[KI77] W. A. Kuperman and F. Ingenito.

Attenuation of the coherent component of sound propagation in shallow water with rough boundaries. J. Acoust. Soc. Am., 61(5): 1178–1187, September 1977.

Kuperman:1980:SCS

[KI80] W. A. Kuperman and F. Ingenito.

Spatial correlation of surface generated noise in a stratified ocean. J. Acoust. Soc. Am., 67(6):1988– 1996, June 1980.

Knight:1981:DSP

[KKK81] R. Knight, W. C. Kay, and S. M.

Kay. Digital signal processing for sonar. Proc IEEE, 69(?), November 1981.

Kloeden:1993:NSS

[KP93] P. E. Kloeden and E. Platen. Numerical solution of stochastic differential equations. In Applications of Mathematics 23. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-54062-8, 0-387-54062-8.

Kraus:1987:MML

[Kra87] D. Kraus. Modifikation von Maximum-Likelihood-Schätzern für Wellenparameter unter Berücksichtigung von Umgebungsrauschen. Master’s thesis, Ruhr-Universität Bochum, 1987.

Kraus:1993:AML

[Kra93] D. Kraus. Approximative Maximum-Likelihood-Schätzung und verwandte Verfahren zur Ortung und Signalschätzung mit Sensorgruppen. PhD thesis, Ruhr-Universität Bochum, D-44780 Bochum, Germany, 1993.

Krolik:1992:MFM

[Kro92] J. L. Krolik. Matched-field minimum variance beamforming in a random ocean channel. J. Acoust. Soc. Am., 92(3):1408– 1419, September 1992.

Kuperman:1985:TAR

[KWG85] W. A. Kuperman, M. F. Werby, and K. E. Gilbert. Towed array response to ship noise: A near-field propagation. In H. G. Urban, editor, Adaptive Methods in Underwater Acoustics, volume 151 of Series C: Mathematical and Physical Sciences, pages 71–80. Kluwer Academic Publisher, Norwell, MA, USA, and Dordrecht, The Netherlands, 1985.

Laval:1985:AMP

[LD85] R. Laval and J. M. Drezet. An analytical method to predict the statistical characteristics of ambient shipping noise. In H. G. Urban, editor, Adaptive Methods in Underwater Acoustics, volume 151 of Series C: Mathematical and Physical Sciences, pages 37–48.

Kluwer Academic Publisher, Norwell, MA, USA, and Dordrecht, The Netherlands, 1985.

Lawson:1974:SLS

[LH74] C. L. Lawson and R. J. Hanson.

Solving Least Squares Problems. Series in Automatic Computation. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1974. ISBN 0-13-822585-0. xii + 340 pp. LCCN QA275 .L425 1974.

Landau:1952:H

[LL52] L. D. Landau and E. M. Lifschitz. Hydrodynamik. Number 6 in Lehrbuch der theoretischen Physik. Verlag Harri Deutsch, Thun, 5. (1991) edition, 1952.

Landau:1953:E

[LL53] L. D. Landau and E. M. Lifschitz. Elastizitätstheorie. Number 7 in Lehrbuch der theoretischen Physik. Verlag Harri Deutsch, Thun, 6. (1989) edition, 1953.

Lewis:1971:ELM

[LO71] T. O. Lewis and P. L. Odell.

Estimation in Linear Models. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1971. ISBN 0-13-289967-1. xi + 193 pp. LCCN QA276.8 .L48.

Lim:1994:UAFa

[LO94a] P. H. Lim and J. M. Ozard. On the underwater acoustic field of a moving point source. part I: Range-independent environment. J. Acoust. Soc. Am., 95(1):131– 137, January 1994.

Lim:1994:UAFb

[LO94b] P. H. Lim and J. M. Ozard. On the underwater acoustic field of a moving point source. part II: Range-dependent environment. J. Acoust. Soc. Am., 95(1):138–151, January 1994.

Li:1990:SLM

[LS90] Shaolin Li and Peter M. Schultheiss. Source localization in a multipath environment using random or deterministic signals. In 5th ASSP Workshop on Spectrum Estimation and Modelling. Rochester, October 1990.

Li:1993:DMR

[LS93] Shaolin Li and Peter M. Schultheiss. Depth measurement of remote sources using multipath propagation. IEEE Journal of Oceanic Engineering, 18(4):379–387, October 1993. Special Issue on Sonar System Technology.

Li:19xx:EUS

[LTxx] S. Li and F. B. Tuteur. Estimation of underwater source parameters by use of multipath information. Fourth IEEE Workshop, 19xx.

Ma:1974:TAA

[Ma74] M. T. Ma. Theory and Application of Antenna Arrays. John Wiley & Sons, New York, NY, USA, 1974. ISBN 0-471-55795-1. xiii + 413 pp. LCCN TK7871.6 .M3.

Maiwald:1989:UKQ

[Mai89] D. Maiwald. Untersuchung eines Kleinste-Quadrate- bzw. eines approximativen Maximum-Likelihood-Zweistufenschätzers zur Analyse stationärer Wellenfelder. Master’s thesis, Ruhr-Universität Bochum, 1989.

Maiwald:1995:BSO

[Mai95] D. Maiwald. Breitbandverfahren zur Signalentdeckung und –ortung mit Sensorgruppen in Seismik– und Sonaranwendungen. PhD thesis, Ruhr-Universität Bochum, May 1995.

Martin:1972:RES

[Mar72] R. D. Martin. Robust estimation of signal amplitude. IEEE Trans. on Information Theory, IT-18: 596–606, 1972.

Maiwald:1992:WPE

[MB92] D. Maiwald and J. F. Böhme.

Wave parameter estimation for transient data. In J. M. F. Moura and I. M. G. Lourtie, editors, Acoustic Signal Processing for Ocean Exploration, NATO ASI Series, pages 321–326. D.Reidel, Dordrecht, 1992.

Maiwald:1994:MTS

[MB94a] D. Maiwald and J. F. Böhme.

Multiple testing for seismic data using bootstrap. In Proc. IEEE Int. Conf. Acoust., Speech, Signal Processing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

Mecklenbraeuker:1994:MFPb

[MB94b] C. F. Mecklenbräuker and J. F.

Böhme. Matched field processing in shallow ocean: Identification of multimode propagation. In L. Bjørnø, editor, Second European Conference on Underwater Acoustics, pages 611–616. European Commission, Copenhagen, July 4-8 1994.

Mecklenbraeuker:1994:MFPa

[MB94c] C. F. Mecklenbräuker and J. F. Böhme. Matched field processing in shallow ocean: Signal arrival identification using EM algorithm. In Proc. IEEE ICASSP, pages II–337–II–340. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 19-23 1994.

Morse:1953:MTP

[MF53] P. M. Morse and H. Feshbach. Methods of Theoretical Physics, Part I and II, Section 2.3, “Motion of Fluids”. McGraw-Hill, New York, NY, USA, 1953.

Mecklenbraeuker:1995:MEE

[MGB95] C. F. Mecklenbräuker, A. Gershman, and J. F. Böhme. MLestimation of environmental parameters in shallow ocean using unknown broadband sources. In Proc. IEEE ICNNSP’95, page to be published. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 1995.

Mecklenbraeuker:1996:BMF

[MGB96] C. F. Mecklenbräuker, M. Geravanchizadeh, and J. F. Böhme. Broadband matched field processing using robust prewhitening and multiple window techniques. In Proc. IEEE ICASSP, page to be published. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1996.

Maiwald:1991:WLS

[MKBB91] D. Maiwald, D. Kraus, G. Bugla, and J. F. Böhme. A weighted least squares criterion for estimation of signal parameters in wavefields. In Int. Conf. on Sonar Signal Processing, pages 95–101. Loughborough, UK, 1991.

Martin-Lauzer:1994:PRI

[MLMS94] F. R. Martin-Lauzer, D. Mauuary, and Y. Stephan. Probabilistic ray identification: A new tool for ocean acoustic tomography. In Proc. IEEE ICASSP, pages II–305–II–308. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 19-23 1994.

Maiwald:1993:MIV

[MMB93a] D. Maiwald, J. W. Dalle Molle, and J. F. Böhme. Model identification and validation of nonstationary seismic signals. In Proc. IEEE Signal Processing Workshop on Higher - Order Statistics. South Lake Tahoe, California, USA, 1993.

Maiwald:1993:PMN

[MMB93b] D. Maiwald, J. W. Dalle Molle, and J. F. Böhme. Parametric methods for nonstationary seismic data analysis. In Proc. SPIE Int. Symp. on Optics, Imaging, and Instrumentation, pages 278–288. San Diego, USA, 1993.

Mecklenbraeuker:1995:FTM

[MMB95] C. F. Mecklenbräuker, D. Maiwald, and J. F. Böhme. F-test in matched field processing: Identifying multimode propagation. In Proc. IEEE ICASSP, pages 3123– 3126. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 8-12 1995.

Milios:1990:ATC

[MN90] E. E. Milios and S. H. Nawab. Acoustic tracking from closest point of approach time, amplitude, and frequency at spatially distributed sensors. J. Acoust. Soc. Am., 87(3):1026–1034, March 1990.

Martin:1982:RRS

[MT82] R. D. Martin and D. J. Thomson. Robust-resistant spectrum estimation. Proceedings of the IEEE, 70 (9):1097–1115, September 1982.

Meyberg:1990:HM

[MV90] Meyberg and Vachenauer. Höhere Mathematik. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990?

Narasimhan:1994:CRB

[NK94] S. Narasimhan and J. L. Krolik. A cramer-rao bound for source range estimation in a random ocean waveguide. In Seventh SP Workshop on Statistical Signal & Array Processing, pages 309–312. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 26-29 1994.

Nehorai:1988:AML

[NL88] A. Nehorai and S. Li. An asymptotic maximum likelihood estimator for source parameters in a multipath environment. 1988.

Nelder:1965:SMF

[NM65] J. A. Nelder and R. Mead. A simplex method for function minimization. The Computer Journal, 7(4):308–313, January 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/hdb/Volume_
07/Issue_04/070308.sgm.abs.
html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
07/Issue_04/tiff/308.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/309.tif;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/
310.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/311.
tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_
07/Issue_04/tiff/312.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_
04/tiff/313.tif.

Nguyen:1994:CPB

[NV94] H. Nguyen and H. L. Van Trees. Comparison of performance bounds for DOA estimation. In Seventh SP Workshop on Statistical Signal & Array Processing, pages 313–316. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 26-29 1994.

Oksendal:1985:SDE

[Øks85] B. Øksendal. Stochastic Differential Equations. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-15292-X. xiii + 205 pp. LCCN QA274.23 .O47 1985.

Burdic:1984:UAS

[Opp84] A. V. Oppenheim, editor. Underwater Acoustic System Analysis. Signal Processing Series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984.

Pekeris:1948:TPE

[Pek48] C. L. Pekeris. Theory of propagation of explosive sound in shallow water. Geol. Soc. Am. Mem., 27, 1948. See also [De 79].

Porat:1983:ESS

[PF83] B. Porat and B. Friedlander. Estimation of spacial and spectral parameters of multiple sources. IEEE Trans. on Information Theory, 29(5):412–428, May 1983.

Park:1989:RSE

[PG89] S. Park and L. A. Gerhardt. A robust spectral estimation by modeling an estimated autocovariance with an ARMA model. IEEE Trans. ASSP, ASSP-37(2): 181, 1989.

Porter:1991:KNM

[Por91] M. B. Porter. The KRAKEN normal mode program. SACLANT-CEN Memorandum SM-245, La Spezia, September 1991.

Porter:1993:AMS

[Por93] Michael B. Porter. Acoustic models and sonar systems. IEEE Journal of Oceanic Engineering, 18(4): 425ff, October 1993.

Preisig:1994:MAA

[Pre94] James C. Preisig. A minimax approach to adaptive matched field processing in an uncertain propagation environment. IEEE Trans. Signal Processing, 42(6): 1305–1316, June 1994.

Priestly:1981:SATa

[Pri81a] M. B. Priestly. Spectral Analysis and Time Series, volume I, Univariate Series of Probability and Mathematical Statistics. Academic Press, New York, NY, USA, 1981.

Priestly:1981:SATb

[Pri81b] M. B. Priestly. Spectral Analysis and Time Series, volume II, Multivariate Series, Prediction and Control of Probability and Mathematical Statistics. Academic Press, New York, NY, USA, 1981.

Purkop:1995:AFT

[Pur95] T. Purkop. Anwendung des F-Tests auf Sonardaten im Flachwasser: Ebene Wellen und Modenausbreitung. Technical report, Ruhr-Universität Bochum, Bochum, Germany, 1995.

Percival:1993:SAP

[PW93] D. B. Percival and A. T. Walden. Spectral Analysis for Physical Applications, Multitaper and Conventional Univariate Techniques. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35532-X (hardcover), 0-521-43541-2 (paperback). xxvii + 583 pp. LCCN QA320 .P434 1993.

Riley:1993:TAE

[RG93] J. L. Riley and D. A. Gray. Towed array estimation using kalman filters – experimental results. IEEE J. Oceanic Eng., 18(4):572–581, October 1993.

Rao:1971:GIM

[RM71] C. R. Rao and S. K. Mitra. Generalized Inverse of Matrices and its Applications. Wiley Series in Probability and Mathematical Statistics. John Wiley & Sons, New York, NY, USA, 1971. ISBN 0-471-70821-6. xiv + 240 pp. LCCN QA263 .R28.

Roach:1970:GF

[Roa70] G. F. Roach. Green’s Functions. Cambridge University Press, Cambridge, UK, 1970. ISBN 0-442-06959-6. xii + 279 pp. LCCN QA372 .R74.

Roach:1982:GF

[Roa82] G. F. Roach. Green’s Functions. Cambridge University Press, Cambridge, UK, second edition, 1982. ISBN 0-521-23890-0 (hardcover), 0-521-28288-8 (paperback). xiii + 325 pp. LCCN QA379 .R6 1982.

Rosenblatt:1985:SSR

[Ros85] M. Rosenblatt. Stationary Sequences and Random Fields. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1985. ISBN 0-8176-3264-6. 258 pp. LCCN QA280 .R67 1985.

Renardy:1993:IPD

[RR93] M. Renardy and R. C. Rogers. An Introduction to Partial Differential Equations, volume 13 of Texts in Applied Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97952-2. xiii + 428 pp. LCCN QA374 .R424 1993.

Rockah:1987:ASC

[RS87] Y. Rockah and P. M. Schultheiss.

Array shape calibration using sources in unknown locations – part I: Far-field sources. IEEE Trans. ASSP, 35(3):286–2xx, March 1987.

Rao:1994:TDA

[RSZ94] C. R. Rao, C. R. Sastry, and B. Zhou. Tracking the direction of arrival of multiple moving targets. IEEE Trans. Signal Processing, 42 (5):1133–1143, May 1994.

Schneider:1995:AMA

[Sch95] H. G. Schneider. Acoustic models at SACLANTCEN. SACLANT-CEN Memorandum SM-285, La Spezia, March 1995.

Skolnik:1962:IRS

[Sko62] M. I. Skolnik. Introduction to Radar Systems. McGraw-Hill, New York, NY, USA, 1962. 648 pp. LCCN TK6575 .S477.

Slepian:1978:PSW

[Sle78] D. Slepian. Prolate spheroidal wave functions, Fourier analysis, and uncertainty- V: The discrete case. The Bell System Technical Journal, 57:1371–1429, 1978. CO-DEN BSTJAN. ISSN 0005-8580.

Schultheiss:1993:OSB

[SM93] P. M. Schultheiss and H. Messer. Optimal and suboptimal broadband source location estimation. IEEE Trans. Signal Processing, 41(9):2752–2763, September 1993. See also [LS90].

Smirnov:19xx:LMB

[Smixx] Smirnov. Lehrbuch der Mathematik, Band IV. 19xx. 228ff pp.

Sobczyk:1985:SWP

[Sob85] K. Sobczyk. Stochastic Wave Propagation. Elsevier Science Publishers B. V., Amsterdam, The Netherlands, 1985. ISBN 0-444-99614-1. viii + 248 pp. LCCN QA927 .S6313 1985. US$90.00.

Sommerfeld:1962:MDM

[Som62a] A. Sommerfeld. Mechanik der deformierbaren Medien. Number 2 in Vorlesungen über Theoretische Physik. Akademische Verlagsgesellschaft, Leipzig, 1962.

Sommerfeld:1962:PDP

[Som62b] A. Sommerfeld. Partielle Differentialgleichungen der Physik. Number 6 in Vorlesungen über Theoretische Physik. Akademische Verlagsgesellschaft, Leipzig, 1962.

Sullivan:1987:PRS

[SR87] E. J. Sullivan and K. Rameau.

Passive ranging with the SNAP model: A performance study. SACLANTCEN Report SR-118, La Spezia, July 1987.

Stergiopoulos:1988:ETA

[SS88] S. Stergiopoulos and E. J. Sullivan. Extended towed-array processor by overlapped correlator. SACLANTCEN Memorandum SM-214, La Spezia, November 1988.

Serebryakov:1995:CEI

[SSM95] G. V. Serebryakov, D. V.

Sidorovitch, and C. F. Mecklenbräuker. Coherence effects of the interference on the performance of optimum/adaptive beamformer. In Proc. IEEE ICASSP, pages 3631–3634. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 8-12 1995.

Stoker:1957:WWM

[Sto57] J. J. Stoker. Water Waves, The Mathematical Teory with Applications. Wiley Classics Library. John Wiley & Sons, New York, NY, USA, reprinted 1992 edition, 1957. ISBN 0-471-57034-6.

Schneider:1985:MSA

[STW85] H. G. Schneider, R. Thiele, and P. C. Wille. Measurement of sound absorption in low salinity water of the baltic sea. J. Acoust. Soc. Am., 77(4):1409–1412, April 1985.

Stoica:1994:OLP

[SVOK94] P. Stoica, M. Viberg, B. Ottersten, and T. Kailath. OPTI-MAL LOCALIZATION OF PAR-TIALLY KNOWN SIGNALS IN UNKNOWN NOISE FIELDS. Proc. IEEE ICASSP, 4:217, April 19-22 1994.

Swingler:1994:SAC

[Swi94] D. N. Swingler. Simple approximation to the cramer rao lower bound on directions of arrival for closely spaced sources. IEEE Trans. ASSP, 41(4):1668– 1672, April 1994.

Tarantola:1986:IPT

[Tar86] A. Tarantola. Inverse Problem Theory. Elsevier Science Publishers B. V., Amsterdam, The Netherlands, 1986. ISBN 0-444-42765-1. xvi + 613 pp. LCCN QA371 .T36 1987.

Turchin:1995:SAA

[TG95] V. L. Turchin and A. B. Gershman. Sensor array approach to nonwave field processing. In Proc. IEEE ICASSP. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 4-8 1995.

Thiel:1987:ACM

[Thi87] R. Thiel. The application of coupled-mode theory to propagation in shallow water with randomly varying sound speed. SACLANTCEN Memorandum SM-199, La Spezia, June 1987.

Thomson:1977:SET

[Tho77] D. J. Thomson. Spectrum estimation techniques for characterization and development of WT 4 waveguide – part I and II. The Bell System Technical Journal, 56: 1769–1815, 1977. CODEN BST-JAN. ISSN 0005-8580.

Thomson:1980:QIS

[Tho80] D. J. Thomson. Quadratic-inverse spectrum estimates: Applications to palaeoclimatology. Phil. Trans. Royal Society London, A(332): 539–597, 1980.

Thomson:1982:SEH

[Tho82] D. J. Thomson. Spectrum estimation and harmonic analysis. Proceedings of the IEEE, 70(9):1055– 1096, September 1982.

Thomson:1994:JMW

[Tho94] D. J. Thomson. Jackknifing multiple-window spectra. In Proc. IEEE ICASSP, pages VI–73–VI– 76. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 19-22 1994.

Urick:1967:PUS

[Uri67] R. J. Urick. Principles of Underwater Sound. McGraw-Hill, New York, NY, USA, second edition, 1967. ISBN 0-07-066085-9. x + 342 pp. LCCN QC225 .U7.

Viberg:1989:SFC

[Vib89] M. Viberg. Subspace Fitting Concepts in Sensor Array Processing. PhD thesis, Linköping University, 1989.

Viberg:1991:EUN

[Vib91] M. Viberg. Effects of unknown noise covariance on parametric array processing algorithms. In Proc. 9th IFAC/IFORS Symp. Identification, System Parameter Estimation, pages 1392–1397. Budapest, 1991.

Viberg:1991:SAP

[VO91] M. Viberg and B. Ottersten. Sensor array processing based on subspace fitting. IEEE Trans. Signal Processing, 39:1110–1121, 1991.

Viberg:1991:DES

[VOK91] M. Viberg, B. Ottersten, and T. Kailath. Detection and estimation in sensor arrays using weighted subspace fitting. IEEE Trans. Signal Processing, 39:2436– 2449, 1991.

Viberg:1994:BAD

[VS94] M. Viberg and A. Swindlehurst. A BAYESIAN APPROACH TO DIRECTION FINDING WITH PARAMETRIC ARRAY UN-CERTAINTY. Proc. IEEE ICASSP, 4:85–88, April 19-22 1994.

Watson:1966:TTB

[Wat66] G. N. Watson. A Treatise on the Theory of Bessel Functions. Cambridge University Press, Cambridge, UK, second edition, 1966. ISBN 0-521-09382-1. vi + 804 pp. LCCN QA 408 W33t 1966.

Wille:1985:ANC

[Wil85] P. C. Wille. Ambient noise: Characteristics of the noise field. In H. G. Urban, editor, Adaptive Methods in Underwater Acoustics, volume 151 of Series C: Mathematical and Physical Sciences, pages 13–35. Kluwer Academic Publisher, Norwell, MA, USA, and Dordrecht, The Netherlands, 1985.

Xu:1994:CWM

[Xu94] Y. Xu. Continuous wave mode structure and constraint beamforming in a waveguide with unknown large inclusions. J. Acoust. Soc. Am., 95(1):152–164, January 1994.

Ziomek:1995:FAF

[Zio95] L. J. Ziomek. Fundamentals of Acoustic Field Theory and Space-Time Signal Processing. CRC Press, 2000 Corporate Blvd., Boca Raton, FL 33431, USA, 1995. ISBN 0-8493-9455-4. xiii + 692 pp. LCCN QC233 .Z66 1995.